OpenStudy (megannicole51):

Split into partial fractions.....((64)/(64-x^2))

OpenStudy (megannicole51):

I get that you have to factor out the denominator which gives you (8-x)(x+8)...but then what?

OpenStudy (loser66):

\[\frac{64}{(8+x)(8-x)}=\frac{A}{(8+x)}+\frac{B}{(8-x)}\] therefore, A (8-x) + B (8+x) = 64 when x =8 \(\rightarrow \)A (8-x) =0 and B (8+8) = 16 B = 64 \(\rightarrow\) B = 4 when x = -8 \(\rightarrow\) immitate my step above to find out A, then plug A and B back to the partial fraction above, you have the answer

OpenStudy (megannicole51):

got it! thank you:) so then how do u finish the integral after that? I have to split it up and then solve it

OpenStudy (megannicole51):

actually i dont have it....im still confused

OpenStudy (anonymous):

you have to multiply to get the common denominator under each term on the right hand side (rhs). after that, you set each part equivalent.

OpenStudy (anonymous):

and solve

OpenStudy (megannicole51):

so to find A i do.....A(8+-8)+B(8+x)=64

OpenStudy (anonymous):

yes... you'll get two equations, one for the number/constant term and one for the x term

OpenStudy (anonymous):

\[8A + 8B = 64 \text{, and } -Ax + Bx = 0x\]

OpenStudy (anonymous):

Once you've found A & B then put them back into the partial fraction decomposition and integrate the terms separately.

OpenStudy (megannicole51):

so would A=4 as well?

OpenStudy (megannicole51):

I just plugged b=4 into 8A+8B=64

OpenStudy (anonymous):

4? I think both A & B are 1...

OpenStudy (megannicole51):

the dude who was helping me before said b=4?

OpenStudy (anonymous):

yeah, sorry. Both A & B are 4.

OpenStudy (megannicole51):

yay okay! so then to split up the equations they would be ((4)/(8-x)) and ((4)/(8+x))?

OpenStudy (anonymous):

so you get \[ \int \frac{64}{64-x^{2}}\, dx =\int \left(\frac{4}{8-x} + \frac{4}{8-x}\right) dx \]

OpenStudy (anonymous):

an you integrate those?

OpenStudy (anonymous):

can...

OpenStudy (megannicole51):

no

OpenStudy (megannicole51):

and shouldnt one of them be 8+x in the numerator?

OpenStudy (anonymous):

no... the decomposition written above is what you want... \[\int \frac{4}{8-x}\, dx + \int \frac{4}{8+x}\, dx \] let's do this first one together...

OpenStudy (megannicole51):

okay:) thank you my notes are so confusing

OpenStudy (anonymous):

\[\int \frac{4}{8-x}\, dx \text{ Let } u=8-x \text{, } du = -dx \text{. Then} \] \[ \int \frac{4}{8-x}\, dx = -4\int \frac{-dx}{8-x} = -4\int \frac{du}{u} = -4\ln u + C = -4\ln(8-x) +C\]

OpenStudy (megannicole51):

oooh u substitution! okay yeah that makes sense!

OpenStudy (anonymous):

it's pretty helpful, that one!

OpenStudy (megannicole51):

wait so how is that the integration of 64/64-x^2?

OpenStudy (anonymous):

that's only the first part... you need to do the second integral, too.

OpenStudy (anonymous):

\[\int \frac{4}{8+x}\, dx\]

OpenStudy (megannicole51):

so would it be the same as the first one except its (8+x)?

OpenStudy (anonymous):

No... no negative. Just practice doing it all until you get comfortable with it all. Otherwise, it's easy to make mistakes and you won't have a solid methodology to fall back on when things get tough or don't work out. Know what I mean? here's a paper on what to do when you have an irreducible quadratic witth 3 terms...

OpenStudy (megannicole51):

oooh okay so it would be 4ln(8+x)+c

OpenStudy (anonymous):

here's a paper that has it all (for partial fraction decomposition and integration)...

OpenStudy (anonymous):

yes!

OpenStudy (megannicole51):

these papers are awesome thank you so much! and so then do you just add each integration?

OpenStudy (anonymous):

take the derivative of that sum and you'll see you get what we integrated, which is equivalent to the original problem statement.

OpenStudy (anonymous):

you're welcome... thanks for being such a great learner!

OpenStudy (megannicole51):

take the derivative of what sum?

OpenStudy (megannicole51):

well thanks for being such a great teacher!

OpenStudy (anonymous):

\[ 4\ln|8+x| - 4\ln|8-x|\]

OpenStudy (anonymous):

I forgot the absolute values for the indefinite integral yielding ln.

OpenStudy (megannicole51):

why is it the absolute value now?

OpenStudy (anonymous):

because of the domain of the ln function. if we had limits of integration, we'd need to make sure whatever we put in the ln function was > 0

OpenStudy (anonymous):

so much to remember!!!

OpenStudy (megannicole51):

so don't they just cancel each other out and ur just left with c?

OpenStudy (megannicole51):

and i know right!!

OpenStudy (anonymous):

no, the arguments are not the same

OpenStudy (anonymous):

\[ \frac{d}{dx} \left( 4\ln|8+x| - 4\ln|8-x| \right) = \frac{4}{8+x} - \frac{-4}{8-x}= \frac{4}{8+x} + \frac{4}{8-x}\]

OpenStudy (megannicole51):

wait so isnt that what we started with? and that makes sense but idk why the absolute values are necessary

OpenStudy (anonymous):

of course it's what we started with! integration is anti-differentiation so it undoes taking the derivative. likewise, once you've integrated, you can take the derivative to check that you get the same thing you just integrated. that let's you know that you've done the integration correctly. As for the absolute values... if x = 9 then ln(8-x) = ln(8-9) = ln(-1) which isn't allowed because the domain of ln y is y>0. here. y = 8-x so 8-x>0 => x<8. Does that make sense? The absolute value is just a precautious technicallity to prevent putting invalid values in the function. the absolute value ensures the argument of the natural log is positive.

OpenStudy (anonymous):

oh, and sorry but it technically should have been this... \[ \frac{d}{dx}\left(4\ln(8+x)-4\ln(x-8)+C\right)=\frac{4}{8+x} + \frac{4}{8-x} \]

OpenStudy (megannicole51):

ooh okay that makes sense. so the answer to the integration of 64/64-x^2 is what u just wrote?

OpenStudy (anonymous):

the derivative of the constant is, of course, 0.

OpenStudy (megannicole51):

and then +c at the end of it?

OpenStudy (anonymous):

yeah... \[ \int \frac{64}{64-x^{2}}\,dx = 4\ln|8+x|-4\ln|8-x|+C\]

OpenStudy (anonymous):

you good?

OpenStudy (megannicole51):

is there another way to write the absolute value? i tried entering it into my online homework stuff and idk how to put the absolute value

OpenStudy (anonymous):

shift \ is an upright bar (|). maybe your online program doesn't require it?

OpenStudy (megannicole51):

i tried just putting parenthesis and it didnt work

OpenStudy (anonymous):

did you try | ?

OpenStudy (megannicole51):

yeah it didnt work:/

OpenStudy (anonymous):

i don't what to say... maybe ask your teacher or someone in your class how to type absolute values in your program. does the program have a help section? it may tell you how to do that.

OpenStudy (megannicole51):

i emailed my professor i guess we will see....thanks for your help!

OpenStudy (anonymous):

you're welcome!